How many grams of K2O are in 1750 mL of 2.5 M solution?
1
Expert's answer
2020-04-27T01:45:41-0400

Solution:

Molarity is the number of moles of solute per liter of solution.

Therefore,

C(K2O) = n(K2O) / V(solution).

Moles of K2O = n(K2O) = C(K2O) × V(solution).

n(K2O) = (2.5 M) × (1.75 L) = 4.375 mol.

n(K2O) = 4.375 mol.


Molar mass is the mass of one mole of a substance.

Molar mass of K2O is 94.196 g/mol.

Therefore,

Mass of K2O = m(K2O) = n(K2O) × M(K2O).

m(K2O) = (4.375 mol) × (94.196 g/mol) = 412.108 g = 412.1 g

m(K2O) = 412.1 g.


Answer: Mass of K2O is 412.1 g.